### 28 Nov 2023 v1

# MVBench: A Comprehensive Multi-modal Video Understanding Benchmark

Abstract
With the rapid development of Multi-modal Large Language Models (MLLMs), a number of diagnostic benchmarks have recently emerged to evaluate the comprehension capabilities of these models. However, most benchmarks predominantly assess spatial understanding in the static image tasks, while overlooking temporal understanding in the dynamic video tasks. To alleviate this issue, we introduce a comprehensive Multi-modal Video understanding Benchmark, namely MVBench, which covers 20 challenging video tasks that cannot be effectively solved with a single frame. Specifically, we first introduce a novel static-to-dynamic method to define these temporal-related tasks. By transforming various static tasks into dynamic ones, we enable the systematic generation of video tasks that require a broad spectrum of temporal skills, ranging from perception to cognition. Then, guided by the task definition, we automatically convert public video annotations into multiple-choice QA to evaluate each task. On one hand, such a distinct paradigm allows us to build MVBench efficiently, without much manual intervention. On the other hand, it guarantees evaluation fairness with ground-truth video annotations, avoiding the biased scoring of LLMs. Moreover, we further develop a robust video MLLM baseline, i.e., VideoChat2, by progressive multi-modal training with diverse instruction-tuning data. The extensive results on our MVBench reveal that, the existing MLLMs are far from satisfactory in temporal understanding, while our VideoChat2 largely surpasses these leading models by over 15% on MVBench. All models and data are available at https://github.com/OpenGVLab/Ask-Anything.
1 Introduction
In the past few years, †† Interns at Shanghai AI Laboratory. Corresponding authors. Multi-modal Large Language Models (MLLMs) [55, 1, 38, 16, 97, 45, 26, 40] have gradually driven the advance in vision-language learning, by plugging visual encoders within various pretrained LLMs [15, 10, 54, 68, 67]. With such a fast development, there is a natural question: How can we evaluate the comprehension capabilities of these MLLMs? Such assessment is vital to confirm their design effectiveness and further improve them for a broader understanding of open-world multi-modalities.
In response to this need, a number of benchmarks have been launched [43, 84, 17, 47, 91], by evaluating MLLMs with Question Answering (QA) formulation of various perception tasks. However, most of these benchmarks primarily concentrate on image-based understanding, where all the questions are designed for spatial perception in the static images, e.g., “Is the man on the stage?”, as shown in Fig. 1. Hence, they suffer from difficulty in assessing temporal evolution in dynamic videos, which is critical to understanding the procedural activities in our realistic world. Recently, several attempts have tried to evaluate MLLMs on temporal perception in videos [36, 81, 57, 49]. But they either work on the very basic video tasks (e.g., action recognition and prediction in SEED-Bench [36]), or focus on the particular domains (e.g., surprising comprehension in FunQA [81]) and restricted scenes (e.g., indoor scenes in Perception Test [57]). As a result, it is limited to leverage these benchmarks to make a comprehensive evaluation on the temporal understanding skills of MLLMs. Besides, they are collected with labor-intensive annotations, leading to expensive manual intervention. To tackle these problems, we propose a Multi-modal Video understanding Benchmark (MVBench), which aims at comprehensively evaluating the temporal perception capabilities of MLLMs in the open world. Compared to these existing benchmarks above, there are two distinct designs in our MVBench.
First, we introduce a novel static-to-dynamic method to systematically define temporal-related tasks, by adapting static image tasks with dynamic evolution. This leads to 20 challenging tasks of video understanding in the MVBench, which covers a wide range of temporal understanding skills from perception to cognition. Specifically, we use static image tasks in the previous multi-modal benchmarks [17, 47] as definition reference. Then, we augment the question of these static tasks with temporal context in the video, e.g., the position task in the image can be flexibly converted into the moving-direction task in the video (“Is the man on the stage?” “What direction is the man moving?”) in Fig. 1. In this case, we can effectively convert all these static tasks into the corresponding dynamic tasks, which cannot be solved without reasoning on the whole video.
Second, guided by the task definition, we design an automatic annotation paradigm to generate multi-choice QAs for each task, by converting 11 public video benchmarks with LLMs. On one hand, it can largely reduce the cost of expensive human annotations. On the other hand, these 11 benchmarks cover various complex domains and diverse scenes, ranging from first-person to third-person perspectives, and from indoor to outdoor environments. Hence, our MVBench is a preferable choice to evaluate the general capability of MLLMs for open-world temporal understanding. More importantly, these benchmarks provide the ground truth for MVBench which guarantees evaluation fairness and accuracy, avoiding biased scoring of LLMs [81, 49].
Finally, we make a thorough evaluation of various well-known MLLMs on our MVBench. Surprisingly, these state-of-the-art image and video MLLMs are far from satisfactory, in terms of temporal perception and cognition. This further motivates us to develop a strong video MLLM baseline, namely VideoChat2, by bridging LLM with a powerful vision foundation model [41]. Subsequently, we introduce a progressive training paradigm with a wide spectrum of multi-modal instructions, allowing effective alignment between video and language. The evaluations show that, our VideoChat2 significantly surpasses the top-performing VideoChat [40] by over 15% accuracy on MVBench, and also achieves the new state-of-the-art results on video conversation [49] and zero-shot QA benchmarks [82, 92]. All the models and data are publicly available, in order to pave the path to general video understanding.
2 Related Works
MLLM. Building upon the significant achievements of Large Language Models (LLMs) [15, 59, 5, 76, 10], scholarly interest has increasingly shifted towards the exploration and development of Multi-modal Large Language Models (MLLMs). This shift aims to augment multi-modal understanding and generation capabilities. Groundbreaking MLLMs such as Flamingo [1] and PaLM-E [16] have seamlessly fused text and vision, setting precedence with their outstanding performances across a range of multi-modal tasks [22, 58, 83, 50]. The recent open-sourcing of LLMs [68, 69, 67, 94, 66] further accelerates the emergence of public MLLMs [97, 45, 20]. Notable examples such as LLaVA [45], MiniGPT-4 [97], and InstructBLIP [11] have contributed by proposing a series of visual instruction-tuning data. Venturing beyond text and static images, several studies have begun harnessing video modality [40, 49, 95, 48], tapping into the vast potential of LLMs for video comprehension tasks [82, 92, 7]. Innovations like VideoChat [40], VideoChatGPT [49], and Valley [48] utilize ChatGPT to generate video instruction-tuning data, aiming to enhance instruction-following capabilities. In the VideoChat2, we aim to critically examine the fundamental temporal understanding capabilities of MLLMs, providing valuable design insights for more robust video MLLMs.
| Spatial | Temporal | Source | Example |
| Action | Action | STAR | What happened after the person took the food? |
| Sequence | (A) Ate the medicine. (B) Tidied up the blanket. (C) Put down the cup/glass/bottle. (D) Took the box. | ||
| Action | STAR | What will the person do next? | |
| Prediction | (A) Put down the pillow. (B) Open the door. (C) Take the book. (D) Open the closet/cabinet. | ||
| Action | PAXION | Which one of these descriptions correctly matches the actions in the video? | |
| Antonym | (A) not sure (B) scattering something down (C) piling something up | ||
| Fine-grained | MiT V1 | What is the action performed by the person in the video? | |
| Action | (A) watering (B) leaking (C) pouring (D) planting | ||
| Unexpected Action | FunQA | What unexpected event contributes to the humor in the video? | |
| (A) The man left without dancing. (B) Two women hugged each other at the end. | |||
| (C) The man finally danced with the woman. (D) Two men hugged each other unexpectedly. | |||
| Object | Object Existence | CLEVRER | Are there any moving green objects when the video ends? (A) not sure (B) yes (C) no |
| Object Interaction | STAR | Which object was tidied up by the person? (A) broom (B) cabinet (C) blanket (D) table | |
| Object Shuffle | Perception Test | Where is the hidden object at the end of the game from the person’s point of view? | |
| (A) Under the first object from the left. (B) Under the third object from the left. | |||
| (C) Under the second object from the left. | |||
| Position | Moving | CLEVRER | What direction is the cyan sphere moving within the video? |
| Direction | (A) The object is stationary. (B) Up and to the right. (C) Down and to the left. (D) Down and to the right. | ||
| Action Localization | Charades- STA | During which part of the video does the action ‘person sitting on a couch’ occur? | |
| (A) In the middle of the video. (B) At the end of the video. | |||
| (C) Throughout the entire video. (D) At the beginning of the video. | |||
| Scene | Scene Transition | MovieNet | What’s the right option for how the scenes in the video change? |
| (A) From the reception desk to the conference room. (B) From the kitchen to the dining area. | |||
| (C) From the server room to the control center. (D) From the classroom to the library. | |||
| Count | Action Count | Perception Test | How many times did the person launch objects on the table? (A) 3 (B) 2 (C) 4 |
| Moving Count | CLEVRER | How many metal objects exit the scene? (A) 2 (B) 3 (C) 1 (D) 0 | |
| Attribute | Moving Attribute | CLEVRER | What shape is the moving object when the video begins? (A) cylinder (B) sphere (C) cube |
| State Change | Perception Test | Is the lighting device on at any point? (A) yes (B) I don’t know (C) no | |
| Pose | Fine-grained Pose | NTU RGB+D | What is the pose performed by the person in the video? (A) pick up (B) sit down (C) drop (D) stand up |
| Character | Character Order | Perception Test | What letter did the person write first on the paper? (A) l (B) v (C) e |
| Cognition | Egocentric Navigation | VLN-CE | For an agent following instruction: “Go left through the door.” What is the next action it should take? |
| (A) Turn left and move forward (B) Move forward (C) Stop (D) Turn right and move forward. | |||
| Episodic Reasoning | TVQA | Why did Castle dress like a fairy when he was speaking to Emily? | |
| (A) To get her to trust him. (B) He secretly loved fairies. (C) He lost a bet with Emily. | |||
| (D) It was dressed like a fairy day at school. (E) Mrs Ruiz made him dress up. | |||
| Counterfactual Inference | CLEVRER | Which of the following will happen if the cylinder is removed? | |
| (A) The cyan rubber object and the blue cube collide. (B) The brown cube collides with the metal cube. | |||
| (C) The cyan rubber object and the metal cube collide. (D) The cyan rubber cube collides with the sphere. | |||
Benchmark. Traditional Vision-Language (VL) benchmarks [30, 21, 83, 80, 82] have primarily honed in on specific capabilities like multi-modal retrieval and vision QA. The rise of MLLMs has catalyzed benchmarks designed for assessing integrated VL tasks. For example, LVLM-eHub [84] provides an interactive model comparison platform through image-related queries. Other benchmarks such as OwlEval [88], MME [17], SEED-Bench [36], MM-Vet [91], and MMBench [47] underscore comprehensive VL skills, introducing evaluation metrics that transcend mere model hierarchies. Meanwhile, the video realm showcased benchmarks like Perception Test [57], examining multi-modal video perception and reasoning, and VideoChatGPT [49] quantifies the capability of dialogue generation from video inputs. FunQA [81] pushes video reasoning limits via counter-intuitive and humorous content. In contrast to the existing benchmarks, MVBench sets itself apart by covering a wide range of temporal tasks, emphasizing temporally-sensitive videos and efficient use of public annotations, and conducting comprehensive evaluations of MLLMs’ temporal understanding.
3 MVBench
In this section, we present our MVBench in detail. We first design the temporal tasks in Tab. 1, and then automatically generate multi-choice QAs for MLLM evaluation in Fig. 2.
3.1 Temporal Task Definition
To design the temporal tasks of MVBench, we introduce a concise static-to-dynamic method by adapting static tasks with dynamic goals. As discussed in the introduction, most existing MLLM benchmarks [17, 47] focus on spatial understanding with systematical definitions of static image tasks. Motivated by this, we propose using these task definitions as references to systematically design temporal tasks, ranging from perception to cognition. As shown in Fig. 1, we begin by summarizing 9 main tasks of spatial understanding from previous benchmarks. Then we enrich these image tasks with video context, creating temporal tasks that can not be effectively solved with a single image and require comprehensive video understanding. Finally, we define 20 temporal tasks as follows. Examples are listed in Tab. 1.
Action. (1) Action Sequence: Retrieve the events occurring before or after a specific action. (2) Action Prediction: Infer the subsequent events based on the current actions. (3) Action Antonym: Distinguish the correct action from two inversely ordered actions. (4) Fine-grained Action: Identify the accurate action from a range of similar options. (5) Unexpected Action: Detect surprising actions in videos characterized by humor, creativity, or magic. Object. (6) Object Existence: Determine the existence of a specific object during a particular event. (7) Object Interaction: Identify the object that participates in a particular event. (8) Object Shuffle: Locate the final position of an object in an occlusion game. Position. (9) Moving Direction: Ascertain the trajectory of a specific object’s movement. (10) Action Localization: Determine the time period when a certain action occurs. Scene. (11) Scene transition: Determine how the scene transitions in the video. Count. (12) Action Count: Calculate how many times a specific action has been performed. (13) Moving Count: Calculate how many objects have performed a certain action. Pose. (14) Fine-grained Pose: Identify the accurate pose category from a range of similar options. Attribute. (15) Moving Attribute: Determine the appearance of a specific moving object at a given moment. (16) State Change: Determine whether the state of a certain object changes throughout the video. Character. (17) Character Order: Determine the order in which the letters appear. Cognition. (18) Egocentric Navigation: Forecast the subsequent action, based on an agent’s current navigation instructions. (19) Episodic Reasoning: Perform reasoning on the characters, events, and objects within an episode of a TV series. (20) Counterfactual Inference: Consider what might happen if a certain event occurs.
3.2 Automatic QA Generation
With the guidance of temporal task definitions, we next collect and annotate videos for each task. Specifically, we design an automatic QA generation paradigm in Fig. 2, which efficiently converts open-sourced video annotations into multi-choice QAs for evaluating MLLMs.
Data Filtration. To reduce the labor-intensive collection, we propose to select videos from existing benchmarks. (1) Video Diversity: To boost video diversity, we carefully select 11 video datasets (see Tab. 1) that cover a broad spectrum of domains and scenes, ranging from first-person to third-person perspectives, and from indoor to outdoor environments. (2) Temporal Sensitivity: To guarantee that each task is temporal sensitive, we eliminate short clips which generally contain negligible motions, and also delete extremely long videos which often present overly complicated contexts that are hard for evaluation. Hence, we select videos with intermediate duration, primarily ranging from 5s to 35s. (3) Question Difficulty: Overly simple or complex questions may lead to indistinguishable evaluations, due to similar responses. To balance the question difficulty, we design the selection criteria for STAR [78] and CLEVRER [29]. For STAR, we enhance the challenge by randomly shifting the start or end points of the video clips, increasing the complexity of localizing specific events. For CLEVRER, we exclude questions that necessitate more than 10 conditions (e.g., material, and shape) for describing specific events, thus decreasing QA difficulty.
QA Generation. Considering that not all the annotations of selected datasets follow the multi-choice QA format, we automatically convert the video annotations into this format via LLMs. Specifically, we first use ChatGPT [54] to generate a question for each video, based on the task definition. Then, we create the corresponding answer options as follows. (1) Template-Based Construction: For most questions, we construct the option candidates directly from the ground truth annotations. For example, the candidates for the Action Antonym task contain the correct action, its opposite action, and a not-sure choice. In the case of the Moving Direction task, the option candidates consist of four directions (i.e., up, down, left, right) and the stationary state. (2) LLM-Based Generation: For the Unexpected Action task in particular, we leverage ChatGPT for converting open-ended QAs into multi-choice QA with answer options. Note that, we use the multi-choice format instead of the open-ended one, for evaluation correction and fairness. This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 200 multi-choice QA pairs for each temporal understanding task. More details of QA generation for all the tasks can be found in the appendix.
Answer Option Processing. For all questions, we randomly sample 3 to 5 answer options from the available candidates, and shuffle the option order, to strengthen the evaluation’s robustness. Additionally, to prevent the common issue of answer leakage where longer options tend to be correct, we further use LLM to guarantee that all the answer options of a question are of similar and reasonable lengths.
3.3 Prompt Design for Evaluation
To emphasize the temporal sensitivity of MLLMs, we craft a detailed system prompt for evaluation (see the bottom right of Fig. 2). This prompt encourages MLLMs to carefully scrutinize video content to answer questions, by paying attention to factors such as the actions and poses of persons, and the details and movements of object movements.
Moreover, another significant challenge lies in extracting options from MLLMs’ responses. MMBench [47] attempts to match predictions with multiple option formats. If failed, it resorts to ChatGPT [54] to extract options through an intricate design. However, this way is relatively inefficient, yielding an alignment rate of only 87% with humans. In contrast, our MVBench employs a simple approach that guarantees 100% rate in option extraction. We enclose the options within parentheses in the questions, and use the answer prompt “Best Option: (” to guide MLLMs for option generation. Results in Tab. 9 demonstrate our prompt’s effectiveness on various MLLMs, allowing us to use accuracy as a reliable metric for evaluation.
4 VideoChat2
After building our MVBench, we evaluate a number of popular image and video MLLMs in Tab. 2. Surprisingly, the existing MLLMs are far from satisfactory in temporal understanding. To fill the gap, we develop a robust video MLLM baseline, which is dubbed as VideoChat2.
4.1 Instruction-Tuning Data
Primarily, the suboptimal performance of MLLMs can be attributed to the limited diversity in instruction-tuning data. To address this issue, we introduce the enriched data as shown in Fig. 3, which comprises 1.9M samples from 34 distinct sources. Following [40, 95], we include both image and video data in the instruction set to improve training.
Motivated by MIT [42], we reorganize all data samples in a uniform format, as shown on the bottom right of Fig. 3. There are two keys involved: {‘image’ or ‘video’}, and {‘QA’}. The first key indicates the path to the vision data. The second key represents a list that contains task instruction (‘i’) and question-answer(‘q’-‘a’). Moreover, different from MIT, which requires researchers to write 10 instructions per dataset, we use ChatGPT to create them, according to {dataset description}, {task description}, and {instruction example} at the top right of Fig. 3. Consequently, our whole instruction-tuning data set can be roughly divided into 6 categories as follows:
(1) Conversation aims at enhancing multi-turn conversational capabilities. We collect conversation data from LLaVA [45] and VideoChat [40]. To expand our data, we integrate the caption data from VideoChatGPT [49] into conversation format based on the video IDs. (2) Simple Caption aims to improve basic visual description capabilities. We choose the widely used COCO Caption [44] and WebVid [3], together with first-order video captions from YouCook2 [13]. (3) Detailed Caption aims at enriching the comprehensive capabilities for understanding visual details. We leverage the detailed caption data from MiniGPT-4 [97], LLaVA [45] and VideoChat [40]. We also integrate Paragraph Captioning [32], TextCaps [62], and TextVR [79], which require uniquely comprehending text within images and videos. (4) VQA aims to improve visual question-answering capabilities. We include the basic VQA (VQAv2 [22], GQA [27], TGIF-QA [28] and WebVidQA [85]), knowledge-based VQA (OK-VQA [50], AOK-VQA [60] and ViQuAE [35]), OCR-based VQA (OCR-VQA [52], TextVQA [63], ST-VQA [4] and DocVQA [51]), and egocentric VQA from Ego4D [23]. (5) Reasoning focuses on enhancing diverse reasoning capacities. We use LLaVA-reasoning [45] and CLEVR [29] for spatial reasoning, VisualMRC [65] for reading comprehension, NExT-QA [80] for temporal reasoning, and CLEVRER [89] for spatiotemporal reasoning. (6) Classification aims at boosting robustness to object and action recognition. We sample data from ImageNet [14], COCO-ITM [44], Kinetics-710 [39] and SthSthV2 [21].
4.2 Progressive Multi-Modal Training
Another critical factor in boosting MLLMs is how to effectively bridge the semantic gap between visual and linguistic representation. To tackle this problem, we adopt a progressive multi-modal training paradigm as shown in Fig. 4.
Stage1: Vision-Language Alignment. In the first stage, we aim at aligning vision and text. To balance efficiency and effectiveness, we freeze the visual encoder and train a flexible QFormer [38], which compresses redundant visual tokens into fewer query tokens, and align these queries with text tokens by multi-modal losses, i.e., Vision-Text Contrastive learning (VTC), Vision-Text Matching (VTM), and Vision-grounded Text Generation (VTG). But different from [38], we choose the pretrained UMT-L [41] as our visual encoder, due to its powerful capability of spatial-temporal representation learning. Moreover, we train QFormer with only 15M image captions from CC3M [61] and CC12M [6] but 10M video captions from WebVid-10M [3], in order to enhance video-language modeling.
| Model | LLM | Avg | AS | AP | AA | FA | UA | OE | OI | OS | MD | AL | ST | AC | MC | MA | SC | FP | CO | EN | ER | CI |
| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
| mPLUG-Owl-I [88] | LLaMA-7B | 29.4 | 25.0 | 20.0 | 44.5 | 27.0 | 23.5 | 36.0 | 24.0 | 34.0 | 23.0 | 24.0 | 34.5 | 34.5 | 22.0 | 31.5 | 40.0 | 24.0 | 37.0 | 25.5 | 21.0 | 37.0 |
| LLaMA-Adapter [96] | LLaMA-7B | 31.7 | 23.0 | 28.0 | 51.0 | 30.0 | 33.0 | 53.5 | 32.5 | 33.5 | 25.5 | 21.5 | 30.5 | 29.0 | 22.5 | 41.5 | 39.5 | 25.0 | 31.5 | 22.5 | 28.0 | 32.0 |
| BLIP2 [38] | FlanT5-XL | 31.4 | 24.5 | 29.0 | 33.5 | 17.0 | 42.0 | 51.5 | 26.0 | 31.0 | 25.5 | 26.0 | 32.5 | 25.5 | 30.0 | 40.0 | 42.0 | 27.0 | 30.0 | 26.0 | 37.0 | 31.0 |
| Otter-I [37] | MPT-7B | 33.5 | 34.5 | 32.0 | 39.5 | 30.5 | 38.5 | 48.5 | 44.0 | 29.5 | 19.0 | 25.5 | 55.0 | 20.0 | 32.5 | 28.5 | 39.0 | 28.0 | 27.0 | 32.0 | 29.0 | 36.5 |
| MiniGPT-4 [97] | Vicuna-7B | 18.8 | 16.0 | 18.0 | 26.0 | 21.5 | 16.0 | 29.5 | 25.5 | 13.0 | 11.5 | 12.0 | 9.5 | 32.5 | 15.5 | 8.0 | 34.0 | 26.0 | 29.5 | 19.0 | 9.9 | 3.0 |
| InstructBLIP [11] | Vicuna-7B | 32.5 | 20.0 | 16.5 | 46.0 | 24.5 | 46.0 | 51.0 | 26.0 | 37.5 | 22.0 | 23.0 | 46.5 | 42.5 | 26.5 | 40.5 | 32.0 | 25.5 | 30.0 | 25.5 | 30.5 | 38.0 |
| LLaVA [45] | Vicuna-7B | 36.0 | 28.0 | 39.5 | 63.0 | 30.5 | 39.0 | 53.0 | 41.0 | 41.5 | 23.0 | 20.5 | 45.0 | 34.0 | 20.5 | 38.5 | 47.0 | 25.0 | 36.0 | 27.0 | 26.5 | 42.0 |
| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
| Otter-V [37] | LLaMA-7B | 26.8 | 23.0 | 23.0 | 27.5 | 27.0 | 29.5 | 53.0 | 28.0 | 33.0 | 24.5 | 23.5 | 27.5 | 26.0 | 28.5 | 18.0 | 38.5 | 22.0 | 22.0 | 23.5 | 19.0 | 19.5 |
| mPLUG-Owl-V [88] | LLaMA-7B | 29.7 | 22.0 | 28.0 | 34.0 | 29.0 | 29.0 | 40.5 | 27.0 | 31.5 | 27.0 | 23.0 | 29.0 | 31.5 | 27.0 | 40.0 | 44.0 | 24.0 | 31.0 | 26.0 | 20.5 | 29.5 |
| VideoChatGPT [49] | Vicuna-7B | 32.7 | 23.5 | 26.0 | 62.0 | 22.5 | 26.5 | 54.0 | 28.0 | 40.0 | 23.0 | 20.0 | 31.0 | 30.5 | 25.5 | 39.5 | 48.5 | 29.0 | 33.0 | 29.5 | 26.0 | 35.5 |
| VideoLLaMA [95] | Vicuna-7B | 34.1 | 27.5 | 25.5 | 51.0 | 29.0 | 39.0 | 48.0 | 40.5 | 38.0 | 22.5 | 22.5 | 43.0 | 34.0 | 22.5 | 32.5 | 45.5 | 32.5 | 40.0 | 30.0 | 21.0 | 37.0 |
| VideoChat [40] | Vicuna-7B | 35.5 | 33.5 | 26.5 | 56.0 | 33.5 | 40.5 | 53.0 | 40.5 | 30.0 | 25.5 | 27.0 | 48.5 | 35.0 | 20.5 | 42.5 | 46.0 | 26.5 | 41.0 | 23.5 | 23.5 | 36.0 |
| VideoChat2 | Vicuna-7B | 34.7 | 24.5 | 27.0 | 49.5 | 27.0 | 38.0 | 53.0 | 28.0 | 40.0 | 25.5 | 27.0 | 38.5 | 41.5 | 27.5 | 32.5 | 46.5 | 26.5 | 36.0 | 33.0 | 32.0 | 40.0 |
| VideoChat2 | Vicuna-7B | 51.1 | 66.0 | 47.5 | 83.5 | 49.5 | 60.0 | 58.0 | 71.5 | 42.5 | 23.0 | 23.0 | 88.5 | 39.0 | 42.0 | 58.5 | 44.0 | 49.0 | 36.5 | 35.0 | 40.5 | 65.5 |
| Evaluation Aspect | VideoChat[40] | VideoChatGPT[49] | VideoChat2 |
| Correctness of Information | 2.23 | 2.40 | 3.02 |
| Detail Orientation | 2.50 | 2.52 | 2.88 |
| Contextual Understanding | 2.53 | 2.62 | 3.51 |
| Temporal Understanding | 1.94 | 1.98 | 2.66 |
| Consistency | 2.24 | 2.37 | 2.81 |
| Avg | 2.29 | 2.38 | 2.98 |
| Model | MSVD-QA | MSRVTT-QA | ANet-QA | |||
| Acc | Score | Acc | Score | Acc | Score | |
| VideoLLaMA [95] | 51.6 | 2.5 | 29.6 | 1.8 | 12.4 | 1.1 |
| VideoChat [40] | 56.3 | 2.8 | 45.0 | 2.5 | 26.5 | 2.2 |
| VideoChatGPT [49] | 64.9 | 3.3 | 49.3 | 2.8 | 35.2 | 2.7 |
| VideoChat2 | 70.0 | 3.9 | 54.1 | 3.3 | 49.1 | 3.3 |
Stage2: Vision-Language Connection. After initial alignment, we then connect the visual encoder with the pretrained LLMs, for building vision-language understanding capabilities. Following [38], we apply a linear projection to further transform the query tokens, and concatenate the projected tokens with the text tokens into LLM for vision-based caption generation (i.e., VTG). But different from [38], we unfreeze the visual encoder for better alignment with LLM. In addition to the aforementioned training data in Stage1, we further introduce 2M image captions (COCO [44], Visual Genome [33], and SBU [56]) and 10M video captions (InternVid [74]), to enrich the caption diversity.
Stage3: Instruction Tuning. In the final stage, we employ the proposed data in Section 4.1 for instruction tuning. To better align responses with instructions, we use low-rank adaptation [24] on the frozen LLM, and tune it along with the visual encoder and QFormer by VTG loss. Moreover, inspired by [11], we integrate instructions (i.e., ‘i’ of ‘QA’) into QFormer, in order to extract instruction-relevant visual tokens as input to LLM. However, different from [11], we do not incorporate questions (i.e., ‘q’ of ‘QA’) into QFormer due to the subpar performances (see appendix.).
5 Experiments
Implementation Details. For visual encoder and LLM, we apply UMT-L [41] and Vicuna-7B v0 [67] by default. Following BLIP2 [38], we deploy QFormer using the pretrained BERT [15]. 32 queries are used in Stage1, and extra 64 queries are introduced in Stage2 and Stage3 when the visual encoder is unfrozen. For efficient training, 4-frame videos are processed through 10 epochs in Stage1 and 1 epoch in Stage2. Transitioning to Stage3, we shift to 8-frame videos for 3 epochs. For evaluation, we input 16-frame videos with elaborate prompts for better results. More details can be found in the appendix.
5.1 Results on MVBench
Tab. 2 presents the evaluation results on MVBench, revealing that current image and video MLLMs are underperforming. For instance, VideoChat [40], a top-performing video MLLM, only marginally surpasses VideoChat2 by 0.8% in average accuracy (35.5% vs. 34.7%), with the latter generating responses from text alone. In contrast, our VideoChat2 markedly exceeds the leading model by over 15%, particularly shining in categories like action, object, scene, attribute, and pose recognition. However, it struggles in position, count, and character tasks, performing less effectively than VideoChat2, which could be attributed to the lack of exposure to these tasks during instruction tuning.
5.2 More Comparisons
Following [49], we use ChatGPT [54] to conduct quantitative comparisons among video MLLMs. (1) Video Conversation: Tab. 3 shows the results on the benchmark of [49]. Compared with VideoChatGPT [49], our VideoChat2 exhibits superior performances across all metrics, with distinct advancements in terms of information correctness as well as context and temporal understanding. This indicates that our VideoChat2 is more adept at comprehending both spatial and temporal details and providing consistent and reliable responses. (2) Zero-Shot Video QA: Tab. 4 lists results of typical video QA datasets [82, 92]. It is evident that our VideoChat2 surpasses all other methods, particularly excelling in understanding long videos in ActivityNet [92].
We further present a qualitative comparison in Fig. 5, where VideoChat2 delivers a precise and thorough response. For more qualitative analyses, see the appendix.
| Data Source | Type | Task | #Num | Avg |
| VideoChat [40] | IV | DCRC | 17K | 36.4 |
| VideoChatGPT [49] | V | DC | 100K | 34.3 2.1 |
| Ours | I | ALL | 1.1M | 42.1 5.7 |
| V | ALL | 0.8M | 50.5 14.1 | |
| IV | ALL | 1.1M | 50.7 14.3 | |
| IV | ALL | 1.9M | 51.1 14.7 | |
| Visual Encoder | LLM | LoRA | Avg |
| EVA-CLIP-g [64] | Vicuna-7B v0 | ✗ | 42.4 |
| ✓ | 45.3 2.9 | ||
| UMT-L [41] | Vicuna-7B v0 | ✗ | 48.6 |
| ✓ | 51.1 2.5 | ||
| Vicuna-13B v0 | ✓ | 51.4 | |
| Vicuna-7B v1.5 | ✗ | 48.1 | |
| ✓ | 51.2 3.1 | ||
| Vicuna-13B v1.5 | ✓ | 51.6 | |
| Stage2 | Stage3 | Avg | ||||
| Visual Encoder | QFomer | Visual Encoder | QFomer | |||
[…]
X |
[…]
X |
[…]
X |
[…]
X |
38.5 | ||
[…]
X |
[…]
10cm |
[…]
X |
[…]
10cm |
47.0 8.5 | ||
[…]
10cm |
[…]
10cm |
[…]
X |
[…]
10cm |
47.5 9.0 | ||
[…]
10cm |
[…]
10cm |
[…]
10cm |
[…]
10cm |
51.1 12.6 | ||
X
and
[…]
10cm
refer to freezing and tuning.
We efficiently freeze the visual encoder in Stage1 and LLM in all stages,
while tuning the visual encoder and QFormer in Stage2&3.
5.3 Ablations of VideoChat2
In this section, we conduct comprehensive analyses of the instruction data, model architecture, and prompt designs.
Instruction Data. Tab. 5 demonstrates that the limited instruction data proposed in VideoChat [40] (17K) and VideoChatGPT [49] (100K) is insufficient for temporal understanding. As we increase the data diversity and quantity, the performances are significantly improved, wherein video data contributes more than image data (50.5% vs. 42.1%). Considering the potential redundancy in the simple caption data of COCO [44] and WebVid [3], we randomly compress them. This results in only a minimal impact on performance (50.7% vs. 51.1%), while accelerating the tuning by 1.7.
Architecture. (1) Visual Encoder: In Tab. 6, we first apply EVA-CLIP-g [64] akin to VideoChat, which achieves 6.9% higher accuracy with our instruction data (42.4% vs. 35.5% for original one in Tab. 2). Further substitutions with UMT-L improve the performance by an additional 6.2%, which demonstrates the effectiveness of our visual encoder. (2) LLM: However, incorporating larger and newer LLMs offers a marginal improvement in the results, indicating that MVBench relies predominantly on the visual encoder. Notably, LoRA [24] consistently uplifts the results, potentially due to its enhanced capacity for instruction following.
| System Prompt | Avg |
| Carefully observe the video and choose the best option for the question. | 49.9 |
| Carefully watch the video and pay attention to the cause, sequence of events, and object details and movements. Based on your observations, select the best option that accurately addresses the question. | 50.5 0.6 |
| Carefully watch the video and pay attention to the cause and sequence of events, the detail and movement of objects and the action and pose of persons. Based on your observations, select the best option that accurately addresses the question. | 51.1 1.2 |
| Model | Answer Prompt | Hit Ratio | Avg |
| VideoChat [40] | 78.2% | 22.8 | |
| Best option: ( | 100% | 35.5 12.7 | |
| VideoChatGPT [49] | 64.6% | 22.0 | |
| Best option: ( | 100% | 32.8 10.8 | |
| VideoChat2 | 96.4% | 50.1 | |
| Best option: ( | 100% | 51.1 1.0 | |
Training Method. Initially, we tune only the linear projection while freezing the visual encoder and QFormer as in MiniGPT-4 [97], but it yielded subpar results in Tab. 7. By unfreezing QFormer as [11], we achieve an 8.5% performance boost. Further, when we unfreeze the visual encoder, results consistently improved, emphasizing the value of more learnable parameters for visual adaptation.
Prompt Design. Tab. 8 reveals that a comprehensive system prompt, which underscores the task requirement, enhances task completion effectiveness. Different from the unstable ChatGPT-extracting methods [47] and more time-consuming log-likelihood comparisons [36], we apply a simple yet effective answer prompt to extra the options. Results in Tab. 9 demonstrate that it accurately targets the option and enhances response precision across various MLLMs. More importantly, VideoChat2 follows the instructions better to return options even without the prompt.
6 Conclusion
This paper introduces MVBench, a comprehensive benchmark for evaluating the temporal understanding capabilities of MLLMs. Moreover, we propose a robust video MLLM baseline, VideoChat2, outperforming the leading models by over 15% on MVBench. Our extensive analyses further direct the designs of MLLMs for temporal understanding.
| config | Stage1 | Stage2 | Stage3 |
| input frame | 4 | 4 | 8 |
| input resolution | 224 | 224 | 224 |
| max text length | 32 | 32 | 512 |
| optimizer | AdamW | ||
| optimizer momentum | |||
| weight decay | 0.02 | ||
| learning rate schedule | cosine decay | ||
| learning rate | 1e-4 | 1e-4 | 2e-5 |
| batch size | 2048 | 512 | 128 |
| warmup epochs | 1 | 0.2 | 0.6 |
| total epochs | 10 | 1 | 3 |
| backbone drop path | 0 | ||
| QFormer drop path | 0.2 | ||
| QFormer dropout | 0 | 0.1 | 0.1 |
| QFormer token | 32 | 96 | 96 |
| flip augmentation | yes | ||
| augmentation | MultiScaleCrop [0.5, 1] | ||
Appendix A Training Hyperparameters
The hyperparameters utilized in different stages of training are listed in Tab. 10. We adopt TSN [71] sampling for all the videos as previous methods [41, 72, 40]. For both Stage1 and Stage2, we employ large-scale image and video caption data, as outlined in the main manuscript. During Stage3, we make use of diverse instruction data and incorporate LoRA modules [24] into the LLM with a rank of 16, an alpha value of 32, and a dropout rate of 0.1. We apply flash attention [12] for LLM to expedite the training process.
Appendix B More Ablations
We have carried out further ablation studies, the results of which are displayed in Tabs. 11, 13, 12, and 14.
| #Query | Instruction | Question | Avg | |
| 32 + 0 | ✓ | ✗ | 47.8 | |
| 32 + 32 | ✓ | ✗ | 50.6 2.8 | |
| 32 + 64 | ✓ | ✗ | 51.1 3.3 | |
| 32 + 96 | ✓ | ✗ | 50.7 2.9 | |
| 32 + 64 | ✓ | ✓ | 50.8 3.0 | |
| 32 + 64 | ✗ | ✗ | 50.5 2.7 | |
QFormer. Considering the richer information of video, we further introduce extra random-initialized queries after Stage1. Tab. 11 shows that more queries in Stage2 and Stage3 is beneficial, leading us to adopt 64 queries by default. Furthermore, inserting instructions without a question effectively steers toward more accurate responses. We argue that overly long context (“instruction question”) may be difficult for information extraction of QFormer.
| Training | Testing | Avg |
| 8224224 | 8224224 | 50.6 |
| 8384384 | 49.9 0.7 | |
| 16224224 | 51.1 0.5 | |
| 32224224 | 51.1 0.5 | |
| 64224224 | 51.0 0.4 | |
| 16224224 | 16224224 | 51.0 0.4 |
Resolution & Frame. Tab. 12 reveals that increasing resolution does not improve performance; however, augmenting the number of frames enhances outcomes. This suggests that our MVBench primarily relies on temporal understanding instead of spatial understanding capacity.
Instruction data. Note that there is a minimal source gap between our instruction data and MVBench. Specifically, the CLEVRER [89] in our instruction data has similar questions as Moving Attribute and Counterfactual Inference in MVBench, leading the evaluation is not strictly out-domain. And the videos of Action Antonym are from SthSthV2 [21], while the antonym is from PAXION [75]. We try to remove CLEVRER and SthSthV2 in the instruction data to evaluate their impact. The results outlined in Tab. 13 suggest a more pronounced influence from CLEVRER data, while SthSthV2 data appears to have less effect.
| Question Prompt | Avg |
| Only give the best option. | 51.1 |
| Only give the best option without any explanation. | 50.9 0.2 |
| Let’s think step by step. Only give the best option. | 50.5 0.6 |
Question prompt. During our experiments, we observed that various MLLMs often provide options along with detailed explanations. To circumvent this, we intentionally craft our question prompts to prevent such detailed outputs. Additionally, drawing inspiration from the Chain-of-Thought [77], we introduce the phrase “Let’s think step by step” into our prompts to direct the MLLMs’ reasoning process. However, as indicated by the results in Tab. 14, these tactics appear to have negative consequences.
| Task | Source | Domain | Data Filtration | QA Generation |
| Action Sequence | STAR [78] | Real-world Indoor Third-person | ✓ ✓ ✗ | QA: Directly adopt |
| Action Antonym | PAXION [75] | Real-world&Simulated Indoor&Outdoor Third-person | N/A | Q: ChatGPT generates A: GTAntonym“not sure” |
| Fine-grained Action | MiT V1 [53] | Real-world&Simulated Indoor&Outdoor Third-person | N/A | Q: ChatGPT generates A: Randomly sample 4 actions from top-6 predictions of UMT-L/16 [41] |
| Unexpected Action | FunQA [81] | Real-world Indoor&Outdoor Third-person | ✓ ✓ ✓ | QA: ChatGPT generates from original QA |
| Object Existence | CLEVRER [89] | Simulated Indoor | ✓ ✓ | Q: ChatGPT generates A: “yes”“no”“not sure” |
| Object Interaction | STAR [78] | Real-world Indoor Third-person | ✓ ✓ ✓ | QA: Directly adopt |
| Object Shuflle | Perception Test [57] | Real-world Indoor First&Third-person | ✓ ✓ | QA: Directly adopt |
| Moving Direction | CLEVRER [89] | Simulated Indoor | Select videos where a certain object is either stationary or moving in a single direction | Q: ChatGPT generates A: “stationary” |
| Action Localization | Charades -STA [19] | Real-world Indoor Third-person | ✓ ✓ ✗ | Q: ChatGPT generates A: “start”“end”“middle”“entire” |
| Scene Transition | MovieNet [25] | Real-world Indoor&Outdoor Third-person | Select videos with continuous scene labels | QA: ChatGPT generates from original QA |
| Action Count | Perception Test [57] | Real-world Indoor First&Third-person | ✓ | QA: Directly adopt |
| Moving Count | CLEVRER [89] | Simulated Indoor | ✓ ✓ | Q: ChatGPT generates A: Randomly shift original answer |
| Fine-grained Pose | NTU RGB+D [46] | Real-world Indoor Third-person | Select videos with specific poses | Q: ChatGPT generates A: Randomly select from similar poses |
| Moving Attribute | CLEVRER [89] | Simulated Indoor | ✓ ✓ ✓ ✓ | Q: ChatGPT generates A: Randomly select from candidates |
| State Change | Perception Test [57] | Real-world Indoor First&Third-person | ✓ ✗ | QA: Directly adopt |
| Character Order | Perception Test [57] | Real-world Indoor First&Third-person | ✓ ✓ | QA: Directly adopt |
| Egocentric Navigation | VLN-CE [31] | Simulated Indoor First-person | ✓ ✓ ✓ | Q: ChatGPT generates A:“move forward”“stop” “turn left and move forward” “turn right and move forward” |
| Episodic Reasoning | TVQA [34] | Real-world Indoor&Outdoor Third-person | ✓ | QA: Directly adopt w/o subtitles |
| Counterfactual Inference | CLEVRER [89] | Simulated Indoor | ✓ ✓ | QA: Directly adopt |
Appendix C Details of QA Generation
In Tab. 15, we present a detailed description of our data generation methodology for MVBench. We have designed various strategies based on different data to increase task difficulty and enhance data diversity. For those datasets requiring question generation, we utilize ChatGPT [54] to generate 3 to 5 questions based on the task definitions.
Appendix D Results on Challenging Video QA
| Model | LLM | Avg | AS | AP | AA | FA | UA | OE | OI | OS | MD | AL | ST | AC | MC | MA | SC | FP | CO | EN | ER | CI |
| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
| GPT-4V take 16 frames as input, and the resolution is 512512, while others use small resolution of 224224. | ||||||||||||||||||||||
| GPT-4V [55] | GPT-4 | 49.5 | 80.0 | 40.0 | 30.0 | 50.0 | 60.0 | 60.0 | 90.0 | 40.0 | 20.0 | 60.0 | 100.0 | 40.0 | 20.0 | 50.0 | 40.0 | 70.0 | 50.0 | 20.0 | 60.0 | 10.0 |
| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
| Otter-I [37] | MPT-7B | 35.0 | 50.0 | 30.0 | 20.0 | 30.0 | 40.0 | 50.0 | 50.0 | 30.0 | 30.0 | 10.0 | 40.0 | 40.0 | 20.0 | 40.0 | 40.0 | 40.0 | 30.0 | 30.0 | 40.0 | 40.0 |
| InstructBLIP [11] | Vicuna-7B | 34.0 | 40.0 | 50.0 | 40.0 | 40.0 | 40.0 | 60.0 | 40.0 | 10.0 | 20.0 | 20.0 | 70.0 | 30.0 | 20.0 | 30.0 | 50.0 | 40.0 | 20.0 | 20.0 | 30.0 | 10.0 |
| LLaVA [45] | Vicuna-7B | 34.5 | 40.0 | 30.0 | 20.0 | 30.0 | 60.0 | 50.0 | 30.0 | 20.0 | 30.0 | 30.0 | 40.0 | 20.0 | 40.0 | 30.0 | 40.0 | 30.0 | 40.0 | 20.0 | 50.0 | 40.0 |
| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
| VideoChatGPT [49] | Vicuna-7B | 32.5 | 30.0 | 30.0 | 50.0 | 30.0 | 40.0 | 70.0 | 20.0 | 50.0 | 10.0 | 20.0 | 10.0 | 30.0 | 10.0 | 30.0 | 70.0 | 20.0 | 50.0 | 10.0 | 50.0 | 20.0 |
| VideoLLaMA [95] | Vicuna-7B | 34.0 | 30.0 | 20.0 | 50.0 | 40.0 | 20.0 | 50.0 | 40.0 | 0.0 | 30.0 | 10.0 | 70.0 | 40.0 | 30.0 | 30.0 | 50.0 | 30.0 | 50.0 | 50.0 | 40.0 | 0.0 |
| VideoChat [40] | Vicuna-7B | 36.5 | 50.0 | 10.0 | 50.0 | 20.0 | 50.0 | 60.0 | 70.0 | 40.0 | 40.0 | 40.0 | 50.0 | 30.0 | 20.0 | 40.0 | 50.0 | 20.0 | 30.0 | 10.0 | 50.0 | 0.0 |
| VideoChat2 | Vicuna-7B | 35.0 | 40.0 | 40.0 | 30.0 | 30.0 | 30.0 | 40.0 | 20.0 | 50.0 | 20.0 | 30.0 | 60.0 | 60.0 | 30.0 | 20.0 | 60.0 | 20.0 | 50.0 | 30.0 | 30.0 | 10.0 |
| VideoChat2 | Vicuna-7B | 56.5 | 60.0 | 60.0 | 90.0 | 60.0 | 60.0 | 70.0 | 80.0 | 30.0 | 20.0 | 40.0 | 100.0 | 40.0 | 40.0 | 40.0 | 60.0 | 50.0 | 40.0 | 60.0 | 60.0 | 70.0 |
| Model | Zero-shot | Fine-tuning | ||||||
| Tem. | Cau. | Des. | Avg | Tem. | Cau. | Des. | Avg | |
| All-in-One [70] | - | - | - | - | 48.6 | 48.0 | 63.2 | 50.6 |
| MIST [18] | - | - | - | - | 56.6 | 54.6 | 66.9 | 57.1 |
| HiTeA [87] | - | - | - | - | 58.3 | 62.4 | 75.6 | 63.1 |
| InternVideo [74] | 43.4 | 48.0 | 65.1 | 59.1 | 58.3 | 62.4 | 75.6 | 63.1 |
| SEVILA [90] | 61.3 | 61.5 | 75.6 | 63.6 | 69.4 | 74.2 | 81.3 | 73.8 |
| VideoChat2 | 57.4 | 61.9 | 69.9 | 61.7 | 64.7 | 68.7 | 76.1 | 68.6 |
| Model | STAR | TVQA | ||||
| Int. | Seq. | Pre. | Fea. | Avg | ||
| FrozenBILM [86] | - | - | - | - | - | 29.7 |
| InternVideo [74] | 43.8 | 43.2 | 42.3 | 37.4 | 41.6 | 35.9 |
| SEVILA [90] | 48.3 | 45.0 | 44.4 | 40.8 | 44.6 | 38.2 |
| VideoChat2 | 58.4 | 60.9 | 55.3 | 53.1 | 59.0 | 40.6 |
In Tabs. 17 and 18, we extend the evaluation of our VideoChat2 to other challenging video benchmarks i.e., NExT-QA [80], STAR [78] and TVQA [34]. Different from the previous methods [90], which provide answers by comparing the likelihood of different options, we output the options directly, following the protocol of MVBench. Our results indicate that VideoChat2 not only holds its own against current SOTA methods [73, 90] on NExT-QA but also markedly outperforms them on STAR and TVQA. This underscores the effectiveness and robustness of VideoChat2.
Appendix E Comparisons with GPT-4V
We further conduct evaluations for GPT-4V [55] in Tab. 16. Given the time constraints, we randomly sample 10 multiple-choice QAs for each task. The results indicate that GPT-4V [55] achieved satisfactory performance on our MVBench, demonstrating its considerable capacity for temporal understanding. Notably, our VideoChat2 outperforms it by increasing accuracy by 7%,
10cm
VideoChat2
66.0
2
[…]
14cm
Otter-I
34.5
3
[…]
10cm
VideoChat
33.5
4
[…]
14cm
LLaVA
28.0
5
[…]
10cm
VideoLLaMA
27.5
6
[…]
14cm
mPLUG-Owl-I
25.0
7
[…]
14cm
BLIP2
24.5
8
[…]
10cm
VideoChatGPT
23.5
9
[…]
14cm
LLaMA-Adapter
23.0
10
[…]
14cm
InstructBLIP
20.0
11
[…]
14cm
MiniGPT-4
16.0
10cm
VideoChat2
47.5
2
[…]
14cm
LLaVA
39.5
3
[…]
14cm
Otter-I
32.0
4
[…]
14cm
BLIP2
29.0
5
[…]
14cm
LLaMA-Adapter
28.0
6
[…]
10cm
VideoChat
26.5
7
[…]
10cm
VideoChatGPT
26.0
8
[…]
10cm
VideoLLaMA
25.5
9
[…]
14cm
mPLUG-Owl-I
20.0
10
[…]
14cm
MiniGPT-4
18.0
11
[…]
14cm
InstructBLIP
16.5
10cm
VideoChat2
83.5
2
[…]
14cm
LLaVA
63.0
3
[…]
10cm
VideoChatGPT
62.0
4
[…]
10cm
VideoChat
56.0
5
[…]
14cm
LLaMA-Adapter
51.0
6
[…]
10cm
VideoLLaMA
51.0
7
[…]
14cm
InstructBLIP
46.0
8
[…]
14cm
mPLUG-Owl-I
44.5
9
[…]
14cm
Otter-I
39.5
10
[…]
14cm
BLIP2
33.5
11
[…]
14cm
MiniGPT-4
26.0
10cm
VideoChat2
49.5
2
[…]
10cm
VideoChat
33.5
3
[…]
14cm
Otter-I
30.5
4
[…]
14cm
LLaVA
30.5
5
[…]
14cm
LLaMA-Adapter
30.0
6
[…]
10cm
VideoLLaMA
29.0
7
[…]
14cm
mPLUG-Owl-I
27.0
8
[…]
14cm
InstructBLIP
24.5
9
[…]
10cm
VideoChatGPT
22.5
10
[…]
14cm
MiniGPT-4
21.5
11
[…]
14cm
BLIP2
17.0
10cm
VideoChat2
60.0
2
[…]
14cm
InstructBLIP
46.0
3
[…]
14cm
BLIP2
42.0
4
[…]
10cm
VideoChat
40.5
5
[…]
14cm
LLaVA
39.0
6
[…]
10cm
VideoLLaMA
39.0
7
[…]
14cm
Otter-I
38.5
8
[…]
14cm
LLaMA-Adapter
33.0
9
[…]
10cm
VideoChatGPT
26.5
10
[…]
14cm
mPLUG-Owl-I
23.5
11
[…]
14cm
MiniGPT-4
16.0
10cm
VideoChat2
58.0
2
[…]
10cm
VideoChatGPT
54.0
3
[…]
14cm
LLaMA-Adapter
53.5
4
[…]
14cm
LLaVA
53.0
5
[…]
10cm
VideoChat
53.0
6
[…]
14cm
BLIP2
51.5
7
[…]
14cm
InstructBLIP
51.0
8
[…]
14cm
Otter-I
48.5
9
[…]
10cm
VideoLLaMA
48.0
10
[…]
14cm
mPLUG-Owl-I
36.0
11
[…]
14cm
MiniGPT-4
29.5
10cm
VideoChat2
71.5
2
[…]
14cm
Otter-I
44.0
3
[…]
14cm
LLaVA
41.0
4
[…]
10cm
VideoLLaMA
40.5
5
[…]
10cm
VideoChat
40.5
6
[…]
14cm
LLaMA-Adapter
32.5
7
[…]
10cm
VideoChatGPT
28.0
8
[…]
14cm
BLIP2
26.0
9
[…]
14cm
InstructBLIP
26.0
10
[…]
14cm
MiniGPT-4
25.5
11
[…]
14cm
mPLUG-Owl-I
24.0
10cm
VideoChat2
42.5
2
[…]
14cm
LLaVA
41.5
3
[…]
10cm
VideoChatGPT
40.0
4
[…]
10cm
VideoLLaMA
38.0
5
[…]
14cm
InstructBLIP
37.5
6
[…]
14cm
mPLUG-Owl-I
34.0
7
[…]
14cm
LLaMA-Adapter
33.5
8
[…]
14cm
BLIP2
31.0
9
[…]
10cm
VideoChat
30.0
10
[…]
14cm
Otter-I
29.5
11
[…]
14cm
MiniGPT-4
13.0
14cm
LLaMA-Adapter
25.5
2
[…]
14cm
BLIP2
25.5
3
[…]
10cm
VideoChat
25.5
4
[…]
10cm
VideoChat2
23.0
5
[…]
10cm
VideoChatGPT
23.0
6
[…]
14cm
mPLUG-Owl-I
23.0
7
[…]
14cm
LLaVA
23.0
8
[…]
10cm
VideoLLaMA
22.5
9
[…]
14cm
InstructBLIP
22.0
10
[…]
14cm
Otter-I
19.0
11
[…]
14cm
MiniGPT-4
11.5
10cm
VideoChat
27.0
2
[…]
14cm
BLIP2
26.0
3
[…]
14cm
Otter-I
25.5
4
[…]
14cm
mPLUG-Owl-I
24.0
5
[…]
10cm
VideoChat2
23.0
6
[…]
14cm
InstructBLIP
23.0
7
[…]
10cm
VideoLLaMA
22.5
8
[…]
14cm
LLaMA-Adapter
21.5
9
[…]
14cm
LLaVA
20.5
10
[…]
10cm
VideoChatGPT
20.0
11
[…]
14cm
MiniGPT-4
12.0
10cm
VideoChat2
88.5
2
[…]
14cm
Otter-I
55.0
3
[…]
10cm
VideoChat
48.5
4
[…]
14cm
InstructBLIP
46.5
5
[…]
14cm
LLaVA
45.0
6
[…]
10cm
VideoLLaMA
43.0
7
[…]
14cm
mPLUG-Owl-I
34.5
8
[…]
14cm
BLIP2
32.5
9
[…]
10cm
VideoChatGPT
31.0
10
[…]
14cm
LLaMA-Adapter
30.5
11
[…]
14cm
MiniGPT-4
9.5
14cm
InstructBLIP
42.5
2
[…]
10cm
VideoChat2
39.0
3
[…]
10cm
VideoChat
35.0
4
[…]
14cm
mPLUG-Owl-I
34.5
5
[…]
14cm
LLaVA
34.0
6
[…]
10cm
VideoLLaMA
34.0
7
[…]
14cm
MiniGPT-4
32.5
8
[…]
10cm
VideoChatGPT
30.5
9
[…]
14cm
LLaMA-Adapter
29.0
10
[…]
14cm
BLIP2
25.5
11
[…]
14cm
Otter-I
20.0
10cm
VideoChat2
42.0
2
[…]
14cm
Otter-I
32.5
3
[…]
14cm
BLIP2
30.0
4
[…]
14cm
InstructBLIP
26.5
5
[…]
10cm
VideoChatGPT
25.5
6
[…]
10cm
VideoLLaMA
22.5
7
[…]
14cm
LLaMA-Adapter
22.5
8
[…]
14cm
mPLUG-Owl-I
22.0
9
[…]
14cm
LLaVA
20.5
10
[…]
10cm
VideoChat
20.5
11
[…]
14cm
MiniGPT-4
15.5
10cm
VideoChat2
58.5
2
[…]
10cm
VideoChat
42.5
3
[…]
14cm
LLaMA-Adapter
41.5
4
[…]
14cm
InstructBLIP
40.5
5
[…]
14cm
BLIP2
40.0
6
[…]
10cm
VideoChatGPT
39.5
7
[…]
14cm
LLaVA
38.5
8
[…]
10cm
VideoLLaMA
32.5
9
[…]
14cm
mPLUG-Owl-I
31.5
10
[…]
14cm
Otter-I
28.5
11
[…]
14cm
MiniGPT-4
8.0
10cm
VideoChatGPT
48.5
2
[…]
14cm
LLaVA
47.0
3
[…]
10cm
VideoChat
46.0
4
[…]
10cm
VideoLLaMA
45.5
5
[…]
10cm
VideoChat2
44.0
6
[…]
14cm
BLIP2
42.0
7
[…]
14cm
mPLUG-Owl-I
40.0
8
[…]
14cm
LLaMA-Adapter
39.5
9
[…]
14cm
Otter-I
39.0
10
[…]
14cm
MiniGPT-4
34.0
11
[…]
14cm
InstructBLIP
32.0
10cm
VideoChat2
49.0
2
[…]
10cm
VideoLLaMA
32.5
3
[…]
10cm
VideoChatGPT
29.0
4
[…]
14cm
Otter-I
28.0
5
[…]
14cm
BLIP2
27.0
6
[…]
10cm
VideoChat
26.5
7
[…]
14cm
MiniGPT-4
26.0
8
[…]
14cm
InstructBLIP
25.5
9
[…]
14cm
LLaMA-Adapter
25.0
10
[…]
14cm
LLaVA
25.0
11
[…]
14cm
mPLUG-Owl-I
24.0
10cm
VideoChat
41.0
2
[…]
10cm
VideoLLaMA
40.0
3
[…]
14cm
mPLUG-Owl-I
37.0
4
[…]
10cm
VideoChat2
36.5
5
[…]
14cm
LLaVA
36.0
6
[…]
10cm
VideoChatGPT
33.0
7
[…]
14cm
LLaMA-Adapter
31.5
8
[…]
14cm
BLIP2
30.0
9
[…]
14cm
InstructBLIP
30.0
10
[…]
14cm
MiniGPT-4
29.5
11
[…]
14cm
Otter-I
27.0
10cm
VideoChat2
35.0
2
[…]
14cm
Otter-I
32.0
3
[…]
10cm
VideoLLaMA
30.0
4
[…]
10cm
VideoChatGPT
29.5
5
[…]
14cm
LLaVA
27.0
6
[…]
14cm
BLIP2
26.0
7
[…]
14cm
mPLUG-Owl-I
25.5
8
[…]
14cm
InstructBLIP
25.5
9
[…]
10cm
VideoChat
23.5
10
[…]
14cm
LLaMA-Adapter
22.5
11
[…]
14cm
MiniGPT-4
19.0
10cm
VideoChat2
40.5
2
[…]
14cm
BLIP2
37.0
3
[…]
14cm
InstructBLIP
30.5
4
[…]
14cm
Otter-I
29.0
5
[…]
14cm
LLaMA-Adapter
28.0
6
[…]
14cm
LLaVA
26.5
7
[…]
10cm
VideoChatGPT
26.0
8
[…]
10cm
VideoChat
23.5
9
[…]
14cm
mPLUG-Owl-I
21.0
10
[…]
10cm
VideoLLaMA
21.0
11
[…]
14cm
MiniGPT-4
9.9
10cm
VideoChat2
65.5
2
[…]
14cm
LLaVA
42.0
3
[…]
14cm
InstructBLIP
38.0
4
[…]
14cm
mPLUG-Owl-I
37.0
5
[…]
10cm
VideoLLaMA
37.0
6
[…]
14cm
Otter-I
36.5
7
[…]
10cm
VideoChat
36.0
8
[…]
10cm
VideoChatGPT
35.5
9
[…]
14cm
LLaMA-Adapter
32.0
10
[…]
14cm
BLIP2
31.0
11
[…]
14cm
MiniGPT-4
3.0
Appendix F Leaderboards and Analyses
To facilitate a clear comparison of different MLLMs, we present the leaderboards for different tasks on MVBench in Tab. 19. Overall, our VideoChat2 achieves the highest rank across 15 tasks.
Action & Pose. For tasks associated with action and pose , our VideoChat2 and VideoChat [40] tends to outperform VideoChatGPT [49], underscoring the significance of elaborate video backbones [39, 41] for effective action and pose recognition.
Object & Attribute. In object-related tasks , the performance of image MLLM, i.e. LLaVA [45], compares favorably with our VideoChat2. It could be attributed to its potent attribute recognition capabilities, as illustrated in . Note that VideoChatGPT [49] is tuned from LLaVA, thus achieving similar results on these tasks.
Position & Count & Character. In position-related tasks , none of the models achieve satisfactory results, their performances being analogous to random guessing. For counting and character-related tasks , our VideoChat2 performs similarly and even worse than VideoChat2 without videos (as in Tab. 2). We hypothesize that current MLLMs have difficulty generalizing to localization and counting tasks in the absence of related tuning data. Some recent studies [9, 2, 8] incorporate grounding data and tune the LLM to enhance localizing and discriminating abilities. In our future work, we will explore improvements in VideoChat2’s grounding ability.
Scene. As presented in Tab. 19, our VideoChat2 excels at scene transition tasks, significantly outperforming other models. This showcases its sensitivity to background changes, making it effective in recognizing camera movements as shown in Fig. 7.
Cognition. In cognition tasks , our VideoChat2 encounters difficulties with complex egocentric navigation and episode reasoning. Given the results from FrozenBiLM [86], where the performance for TVQA reasoning significantly improves with the incorporation of speech subtitles, we suggest that visual information alone may not be sufficient. The inclusion of other modalities, such as depth and audio, could prove beneficial.
Appendix G Qualitative Results
Additional qualitative results can be found in Figs. 6 and 7. Compared with VideoChat [40] and VideoChatGPT [49], our VideoChat2 performs admirably across a range of tasks in MVBench. It possesses the capacity to accurately identify the properties of moving objects, recognize unforeseen actions, and predict future movements based on video context. Moreover, it exhibits robustness when dealing with both real and generated videos, adeptly providing detailed insights into human actions, camera motions, background ambiance, and character attributes.

### 21 Jan 2024 v3

@@
 In this case, we can effectively convert all these static tasks into the corresponding dynamic tasks, which cannot be solved without reasoning on the whole video.
-Second, guided by the task definition, we design an automatic annotation paradigm to generate multi-choice QAs for each task, by converting 11 public video benchmarks with LLMs.
+Second, guided by the task definition, we design an automatic annotation paradigm to generate multiple-choice QAs for each task, by converting 11 public video benchmarks with LLMs.
 On one hand, it can largely reduce the cost of expensive human annotations.
@@
 3 MVBench
-In this section, we present our MVBench in detail. We first design the temporal tasks in Tab. 1, and then automatically generate multi-choice QAs for MLLM evaluation in Fig. 2.
+In this section, we present our MVBench in detail. We first design the temporal tasks in Tab. 1, and then automatically generate multiple-choice QAs for evaluation in Fig. 2.
 3.1 Temporal Task Definition
@@
 (13) Moving Count: Calculate how many objects have performed a certain action.
-Pose.
-(14) Fine-grained Pose: Identify the accurate pose category from a range of similar options.
 Attribute.
@@
 (15) State Change: Determine whether the state of a certain object changes throughout the video.
+Pose.
+(16) Fine-grained Pose: Identify the accurate pose category from a range of similar options.
 Character.
@@
 Specifically, we design an automatic QA generation paradigm in Fig.
-2, which efficiently converts open-sourced video annotations into multi-choice QAs for evaluating MLLMs.
+2, which efficiently converts open-sourced video annotations into multiple-choice QAs for evaluating MLLMs.
 Data Filtration.
@@
 QA Generation.
-Considering that not all the annotations of selected datasets follow the multi-choice QA format, we automatically convert the video annotations into this format via LLMs.
+Considering that not all the annotations of selected datasets follow the multiple-choice QA format, we automatically convert the video annotations into this format via LLMs.
 Specifically, we first use ChatGPT [54] to generate a question for each video, based on the task definition.
@@
 In the case of the Moving Direction task, the option candidates consist of four directions (i.e., up, down, left, right) and the stationary state.
-(2) LLM-Based Generation: For the Unexpected Action task in particular, we leverage ChatGPT for converting open-ended QAs into multi-choice QA with answer options.
-Note that, we use the multi-choice format instead of the open-ended one, for evaluation correction and fairness.
+(2) LLM-Based Generation: For the Unexpected Action task in particular, we leverage ChatGPT for converting open-ended QAs into multiple-choice QA with answer options.
+Note that, we use the multiple-choice format instead of the open-ended one, for evaluation correction and fairness.
 This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention.
-Ultimately, we produce 200 multi-choice QA pairs for each temporal understanding task.
+Ultimately, we produce 200 multiple-choice QA pairs for each temporal understanding task.
 More details of QA generation for all the tasks can be found in the appendix.
@@
 | Model | LLM | Avg | AS | AP | AA | FA | UA | OE | OI | OS | MD | AL | ST | AC | MC | MA | SC | FP | CO | EN | ER | CI |
-| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
-| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
+| Random | - | 27.3 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
+| Image MLLMs: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
 | mPLUG-Owl-I [88] | LLaMA-7B | 29.4 | 25.0 | 20.0 | 44.5 | 27.0 | 23.5 | 36.0 | 24.0 | 34.0 | 23.0 | 24.0 | 34.5 | 34.5 | 22.0 | 31.5 | 40.0 | 24.0 | 37.0 | 25.5 | 21.0 | 37.0 |
@@
 | LLaVA [45] | Vicuna-7B | 36.0 | 28.0 | 39.5 | 63.0 | 30.5 | 39.0 | 53.0 | 41.0 | 41.5 | 23.0 | 20.5 | 45.0 | 34.0 | 20.5 | 38.5 | 47.0 | 25.0 | 36.0 | 27.0 | 26.5 | 42.0 |
-| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
+| Video MLLMs: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
 | Otter-V [37] | LLaMA-7B | 26.8 | 23.0 | 23.0 | 27.5 | 27.0 | 29.5 | 53.0 | 28.0 | 33.0 | 24.5 | 23.5 | 27.5 | 26.0 | 28.5 | 18.0 | 38.5 | 22.0 | 22.0 | 23.5 | 19.0 | 19.5 |
@@
 | VideoChat2 | Vicuna-7B | 51.1 | 66.0 | 47.5 | 83.5 | 49.5 | 60.0 | 58.0 | 71.5 | 42.5 | 23.0 | 23.0 | 88.5 | 39.0 | 42.0 | 58.5 | 44.0 | 49.0 | 36.5 | 35.0 | 40.5 | 65.5 |
+| GPT-4V take 16 frames as input, and the resolution is 512512, while others use small resolution of 224224. | ||||||||||||||||||||||
+| GPT-4V [55] | GPT-4 | 43.5 | 55.5 | 63.5 | 72.0 | 46.5 | 73.5 | 18.5 | 59.0 | 29.5 | 12.0 | 40.5 | 83.5 | 39.0 | 12.0 | 22.5 | 45.0 | 47.5 | 52.0 | 31.0 | 59.0 | 11.0 |
 | Evaluation Aspect | VideoChat[40] | VideoChatGPT[49] | VideoChat2 |
@@
 However, it struggles in position, count, and character tasks, performing less effectively than VideoChat2, which could be attributed to the lack of exposure to these tasks during instruction tuning.
+Additionally, we have included an evaluation of the powerful GPT-4V [55] on our MVBench.
+The results show that while GPT-4V [55] achieves satisfactory performance, particularly in demonstrating its capacity for temporal understanding, our VideoChat2 still manages to surpass it, increasing accuracy by 7.6%.
+This further underscores the superiority of our model in handling a broader range of tasks.
 5.2 More Comparisons
@@
 Our extensive analyses further direct the designs of MLLMs for temporal understanding.
+Appendix A Training Hyperparameters
+The hyperparameters used in different stages of training are listed in Tab.
+10.
+We adopt TSN [71] sampling for all the videos as previous methods [41, 72, 40].
+For both Stage1 and Stage2, we employ large-scale image and video caption data, as outlined in the main manuscript.
+During Stage3, we make use of diverse instruction data and incorporate LoRA modules [24] into the LLM with a rank of 16, an alpha value of 32, and a dropout rate of 0.1.
+We apply flash attention [12] to expedite the training process.
 | config | Stage1 | Stage2 | Stage3 |
@@
 | augmentation | MultiScaleCrop [0.5, 1] | ||
-Appendix A Training Hyperparameters
-The hyperparameters utilized in different stages of training are listed in Tab.
-10.
-We adopt TSN [71] sampling for all the videos as previous methods [41, 72, 40].
-For both Stage1 and Stage2, we employ large-scale image and video caption data, as outlined in the main manuscript.
-During Stage3, we make use of diverse instruction data and incorporate LoRA modules [24] into the LLM with a rank of 16, an alpha value of 32, and a dropout rate of 0.1.
-We apply flash attention [12] for LLM to expedite the training process.
 Appendix B More Ablations
 We have carried out further ablation studies, the results of which are displayed in Tabs. 11, 13, 12, and 14.
+QFormer.
+Considering the richer information of video, we further introduce extra random-initialized queries after Stage1.
+Tab.
+11 shows that more queries in Stage2 and Stage3 is beneficial, leading us to adopt 64 queries by default.
+Furthermore, inserting instructions without a question effectively steers toward more accurate responses.
+We argue that overly long context (“instruction question”) may be difficult for information extraction of QFormer.
 | #Query | Instruction | Question | Avg | |
@@
 | 32 + 64 | ✗ | ✗ | 50.5 2.7 | |
-QFormer.
-Considering the richer information of video, we further introduce extra random-initialized queries after Stage1.
-Tab.
-11 shows that more queries in Stage2 and Stage3 is beneficial, leading us to adopt 64 queries by default.
-Furthermore, inserting instructions without a question effectively steers toward more accurate responses.
-We argue that overly long context (“instruction question”) may be difficult for information extraction of QFormer.
 | Training | Testing | Avg |
@@
 | 16224224 | 16224224 | 51.0 0.4 |
+| Question Prompt | Avg |
+| Only give the best option. | 51.1 |
+| Only give the best option without any explanation. | 50.9 0.2 |
+| Let’s think step by step. Only give the best option. | 50.5 0.6 |
 Resolution & Frame.
@@
 13 suggest a more pronounced influence from CLEVRER data, while SthSthV2 data appears to have less effect.
-| Question Prompt | Avg |
-| Only give the best option. | 51.1 |
-| Only give the best option without any explanation. | 50.9 0.2 |
-| Let’s think step by step. Only give the best option. | 50.5 0.6 |
-Question prompt.
-During our experiments, we observed that various MLLMs often provide options along with detailed explanations.
-To circumvent this, we intentionally craft our question prompts to prevent such detailed outputs.
-Additionally, drawing inspiration from the Chain-of-Thought [77], we introduce the phrase “Let’s think step by step” into our prompts to direct the MLLMs’ reasoning process.
-However, as indicated by the results in Tab.
-14, these tactics appear to have negative consequences.
 | Task | Source | Domain | Data Filtration | QA Generation |
@@
 | Moving Count | CLEVRER [89] | Simulated Indoor | ✓ ✓ | Q: ChatGPT generates A: Randomly shift original answer |
-| Fine-grained Pose | NTU RGB+D [46] | Real-world Indoor Third-person | Select videos with specific poses | Q: ChatGPT generates A: Randomly select from similar poses |
 | Moving Attribute | CLEVRER [89] | Simulated Indoor | ✓ ✓ ✓ ✓ | Q: ChatGPT generates A: Randomly select from candidates |
 | State Change | Perception Test [57] | Real-world Indoor First&Third-person | ✓ ✗ | QA: Directly adopt |
+| Fine-grained Pose | NTU RGB+D [46] | Real-world Indoor Third-person | Select videos with specific poses | Q: ChatGPT generates A: Randomly select from similar poses |
 | Character Order | Perception Test [57] | Real-world Indoor First&Third-person | ✓ ✓ | QA: Directly adopt |
@@
 | Counterfactual Inference | CLEVRER [89] | Simulated Indoor | ✓ ✓ | QA: Directly adopt |
+Question prompt.
+During our experiments, we observed that various MLLMs often provide options along with detailed explanations.
+To circumvent this, we intentionally craft our question prompts to prevent such detailed outputs.
+Additionally, drawing inspiration from the Chain-of-Thought [77], we introduce the phrase “Let’s think step by step” into our prompts to direct the MLLMs’ reasoning process.
+However, as indicated by the results in Tab.
+14, these tactics appear to have negative consequences.
 Appendix C Details of QA Generation
@@
 Appendix D Results on Challenging Video QA
-| Model | LLM | Avg | AS | AP | AA | FA | UA | OE | OI | OS | MD | AL | ST | AC | MC | MA | SC | FP | CO | EN | ER | CI |
-| Random | - | 27.7 | 25.0 | 25.0 | 33.3 | 25.0 | 25.0 | 33.3 | 25.0 | 33.3 | 25.0 | 25.0 | 25.0 | 33.3 | 33.3 | 33.3 | 33.3 | 25.0 | 33.3 | 25.0 | 20.0 | 30.9 |
-| GPT-4V take 16 frames as input, and the resolution is 512512, while others use small resolution of 224224. | ||||||||||||||||||||||
-| GPT-4V [55] | GPT-4 | 49.5 | 80.0 | 40.0 | 30.0 | 50.0 | 60.0 | 60.0 | 90.0 | 40.0 | 20.0 | 60.0 | 100.0 | 40.0 | 20.0 | 50.0 | 40.0 | 70.0 | 50.0 | 20.0 | 60.0 | 10.0 |
-| ImageLLM: Following [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM. | ||||||||||||||||||||||
-| Otter-I [37] | MPT-7B | 35.0 | 50.0 | 30.0 | 20.0 | 30.0 | 40.0 | 50.0 | 50.0 | 30.0 | 30.0 | 10.0 | 40.0 | 40.0 | 20.0 | 40.0 | 40.0 | 40.0 | 30.0 | 30.0 | 40.0 | 40.0 |
-| InstructBLIP [11] | Vicuna-7B | 34.0 | 40.0 | 50.0 | 40.0 | 40.0 | 40.0 | 60.0 | 40.0 | 10.0 | 20.0 | 20.0 | 70.0 | 30.0 | 20.0 | 30.0 | 50.0 | 40.0 | 20.0 | 20.0 | 30.0 | 10.0 |
-| LLaVA [45] | Vicuna-7B | 34.5 | 40.0 | 30.0 | 20.0 | 30.0 | 60.0 | 50.0 | 30.0 | 20.0 | 30.0 | 30.0 | 40.0 | 20.0 | 40.0 | 30.0 | 40.0 | 30.0 | 40.0 | 20.0 | 50.0 | 40.0 |
-| VideoLLM: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames. | ||||||||||||||||||||||
-| VideoChatGPT [49] | Vicuna-7B | 32.5 | 30.0 | 30.0 | 50.0 | 30.0 | 40.0 | 70.0 | 20.0 | 50.0 | 10.0 | 20.0 | 10.0 | 30.0 | 10.0 | 30.0 | 70.0 | 20.0 | 50.0 | 10.0 | 50.0 | 20.0 |
-| VideoLLaMA [95] | Vicuna-7B | 34.0 | 30.0 | 20.0 | 50.0 | 40.0 | 20.0 | 50.0 | 40.0 | 0.0 | 30.0 | 10.0 | 70.0 | 40.0 | 30.0 | 30.0 | 50.0 | 30.0 | 50.0 | 50.0 | 40.0 | 0.0 |
-| VideoChat [40] | Vicuna-7B | 36.5 | 50.0 | 10.0 | 50.0 | 20.0 | 50.0 | 60.0 | 70.0 | 40.0 | 40.0 | 40.0 | 50.0 | 30.0 | 20.0 | 40.0 | 50.0 | 20.0 | 30.0 | 10.0 | 50.0 | 0.0 |
-| VideoChat2 | Vicuna-7B | 35.0 | 40.0 | 40.0 | 30.0 | 30.0 | 30.0 | 40.0 | 20.0 | 50.0 | 20.0 | 30.0 | 60.0 | 60.0 | 30.0 | 20.0 | 60.0 | 20.0 | 50.0 | 30.0 | 30.0 | 10.0 |
-| VideoChat2 | Vicuna-7B | 56.5 | 60.0 | 60.0 | 90.0 | 60.0 | 60.0 | 70.0 | 80.0 | 30.0 | 20.0 | 40.0 | 100.0 | 40.0 | 40.0 | 40.0 | 60.0 | 50.0 | 40.0 | 60.0 | 60.0 | 70.0 |
 | Model | Zero-shot | Fine-tuning | ||||||
@@
 In Tabs.
-17 and 18, we extend the evaluation of our VideoChat2 to other challenging video benchmarks i.e., NExT-QA [80], STAR [78] and TVQA [34].
+16 and 17, we extend the evaluation of our VideoChat2 to other challenging video benchmarks i.e., NExT-QA [80], STAR [78] and TVQA [34].
 Different from the previous methods [90], which provide answers by comparing the likelihood of different options, we output the options directly, following the protocol of MVBench.
@@
 This underscores the effectiveness and robustness of VideoChat2.
-Appendix E Comparisons with GPT-4V
-We further conduct evaluations for GPT-4V [55] in Tab.
-16.
-Given the time constraints, we randomly sample 10 multiple-choice QAs for each task.
-The results indicate that GPT-4V [55] achieved satisfactory performance on our MVBench, demonstrating its considerable capacity for temporal understanding.
-Notably, our VideoChat2 outperforms it by increasing accuracy by 7%,
 10cm
@@
 10cm
-VideoChat2
-58.5
-2
-[…]
-10cm
-VideoChat
-42.5
-3
-[…]
-14cm
-LLaMA-Adapter
-41.5
-4
-[…]
-14cm
-InstructBLIP
-40.5
-5
-[…]
-14cm
-BLIP2
-40.0
-6
-[…]
-10cm
-VideoChatGPT
-39.5
-7
-[…]
-14cm
-LLaVA
-38.5
-8
-[…]
-10cm
-VideoLLaMA
-32.5
-9
-[…]
-14cm
-mPLUG-Owl-I
-31.5
-10
-[…]
-14cm
-Otter-I
-28.5
-11
-[…]
-14cm
-MiniGPT-4
-8.0
-10cm
 VideoChatGPT
@@
 24.0
+10cm
+VideoChat2
+58.5
+2
+[…]
+10cm
+VideoChat
+42.5
+3
+[…]
+14cm
+LLaMA-Adapter
+41.5
+4
+[…]
+14cm
+InstructBLIP
+40.5
+5
+[…]
+14cm
+BLIP2
+40.0
+6
+[…]
+10cm
+VideoChatGPT
+39.5
+7
+[…]
+14cm
+LLaVA
+38.5
+8
+[…]
+10cm
+VideoLLaMA
+32.5
+9
+[…]
+14cm
+mPLUG-Owl-I
+31.5
+10
+[…]
+14cm
+Otter-I
+28.5
+11
+[…]
+14cm
+MiniGPT-4
+8.0
 10cm
@@
 3.0
-Appendix F Leaderboards and Analyses
-To facilitate a clear comparison of different MLLMs, we present the leaderboards for different tasks on MVBench in Tab. 19. Overall, our VideoChat2 achieves the highest rank across 15 tasks.
+Appendix E Leaderboards and Analyses
+To facilitate a clear comparison of different open-sourced MLLMs, we present the leaderboards for different tasks on MVBench in Tab.
+18.
+Overall, our VideoChat2 achieves the highest rank across 15 tasks.
 Action & Pose.
@@
 As presented in Tab.
-19, our VideoChat2 excels at scene transition tasks, significantly outperforming other models.
+18, our VideoChat2 excels at scene transition tasks, significantly outperforming other models.
 This showcases its sensitivity to background changes, making it effective in recognizing camera movements as shown in Fig.
@@
 The inclusion of other modalities, such as depth and audio, could prove beneficial.
-Appendix G Qualitative Results
+Appendix F Qualitative Results
 Additional qualitative results can be found in Figs.